What is the wavelength of a 2.99 Hz wave?
1
Expert's answer
2019-10-11T09:39:36-0400

Q96338

Solution:

As we know, "\nu=c/\lambda"

where, "\nu= frequency of wave (in Hz.)"

"c=speed of light=2.99*10^8m\/s."

"\lambda=wavelength of wave(in m.)"

Thus, "\lambda=c/\nu"

"=2.99*10^8\/2.99"

"=10^8 m." (Answer)